# Cross-Domain Few-Shot Object Detection via Enhanced Open-Set Object Detector

### v2

@@ 5 Experiments
Implementation Details. The mAP for 1/5/10 shots are reported as in Distill-cdfsod [44]; Besides the vanilla OVD methods that detect novel objects directly, we adopt the "pretrain, finetune, and test" pipeline for all evaluated methods as in Sec. 3.1. In the case of our CD-ViTO, the base DE-ViT model pretrained on COCO is taken, then the $M_{DET}$, $M_{CLS}$, $M_{LIF}$, $M_{IR}$, and $M_{DP}$ are tuned on novel support images $S$ using the loss as in Eq. 1. Finetuning details e.g., hyperparameters, optimizer, learning rate, and epochs will be attached in the supplementary materials.
@@ 5.1 Evaluated Methods
Typical FSOD Methods. Several FSOD methods are included to examine whether these traditional algorithms can handle the huge domain gap posed by CD-FSOD. Concretely, Meta-RCNN [46], TFA [42], FSCE [38], and DeFRCN [33] are introduced.
@@
Open-Set FSOD/OD Methods. The DE-ViT [49] which shows privilege in FSOD benchmarks is studied. In addition to DE-ViT [49] (vision-only method), we also want to evaluate the performance of classical vision-language matching methods on our datasets. Therefore, we adapt another widely known open-vocabulary object detection method Detic [53]. Apart from the vanilla Detic, we also finetune the model on support images which results in the "Detic-FT".
@@ 5.2 Main Results on CD-FSOD
Can open-set models directly address the challenges of CD-FSOD? Unfortunately, despite being designed to recognize any novel category in the open world and the expectation that well-trained open-set models can make direct inferences, open-set detectors still face challenges posed by domain gap issues. Through the results of Detic and DE-ViT, it becomes evident that the answer is no. It is noteworthy that Detic performs inference using category names and the query image, essentially adopting a zero-shot approach. Consequently, the results for 1/5/10 shots remain the same. On the other hand, DE-ViT requires support images to construct prototypes, leading to variations in results for different shots. These findings highlight the nontrivial nature of solving the CD-FSOD by directly applying existing open-set models.
@@
Do different open-set models exhibit distinct performances? Yes. We delve into the performance of DE-ViT and Detic, two models that achieve open-set learning through entirely different approaches. DE-ViT relies solely on vision features, while Detic relies on the relationship between visual features and category text. Upon analyzing the results of Detic-FT and CD-ViTO, a clear trend emerges. CD-ViTO demonstrates a decisive advantage on ArTaxOr, Clipart1k, DIOR, and DeepFish, but falls behind Detic-FT on NEU-DET and UODD. We observe that when the Instance Boundary (IB) degree of the target dataset is slight or moderate, the vision-based CD-ViTO performs better. Conversely, in scenarios where boundaries are significantly indefinable, indicating strong similarity between foreground and background, methods like Detic-FT, which leverages both visual and text information, prove more competent.
How do style, inter-class variance (ICV), and indefinable boundaries (IB) impact the domain gap? We meticulously analyze the influence of style, ICV, and IB on domain gaps based on experimental results and benchmark analysis in Sec. 3.3. 1) Style, often considered strongly related to the domain, has a relatively smaller impact on CD-FSOD task compared to ICV and IB. For instance, DIOR shares relatively similar ICV and IB with COCO but has a different style. Despite this, the performance of typical FSOD methods does not significantly drop on DIOR. For example, DeFRCN achieves 9.3 mAP on both COCO’s testing set and DIOR under the 1-shot setting. Similarly, the cartoon-style Clipart1k, distinct from the photorealistic COCO, also receives relatively high mAPs. 2) ICV can cause domain issues but appears manageable. ArTaxOr, with the same style and similar IB as COCO, but with fine-grained categories (arthropods) compared to the coarse-level semantic concepts in COCO. This difference in ICV leads to a significant drop in performance for typical FSOD methods on ArTaxOr, such as from 9.1 to 3.6 on 1-shot using DeFRCN. However, we observe that this negative effect can be mitigated by utilizing better vision features, as demonstrated by ViTDeT-FT and our CD-ViTO. 3) The change in IB poses the greatest challenge for CD-FSOD. Nearly all types of methods struggle to achieve satisfactory results on NEU-DET and UODD when significant IB exists. Blurred boundaries make it challenging to detect correct bounding boxes, resulting in poor overall performance.
@@ 5.3 Analysis on the Proposed Modules
To demonstrate the effectiveness of our proposed components, including the finetuning (FT), the learnable instance features $M_{LIF}$, the instance reweighting module $M_{IR}$, and the domain prompter $M_{DP}$, we conduct ablation studies by incrementally adding them to the base DE-ViT method. The 10-shot results are presented in Fig. 4(a).
Overall, we observe that: 1) By simply finetuning top heads, the performance is significantly improved. The results reinforce the observed phenomenon and analysis from Tab. 1 that finetuning is crucial in CD-FSOD. 2) The learnable instance features further boost performance clearly, indicating that we do make the features better by optimizing them. 3) Although not as pronounced as in finetuning and $M_{LIF}$, the results on target datasets still demonstrate the effectiveness of our instance reweighting module and domain prompter. Note that we don’t consider finetuning as a baseline method for adapting DE-ViT, given the expectation that open-set detectors are supposed to recognize novel objects without this process. Exploring how the vanilla open-set detector performs on cross-domain FSOD is one of our goals. Additionally, in supplementary materials, we explore how to finetune the model efficiently. Thus, we include finetuning as our component but don’t claim it as a technical contribution.
@@
More Discussions on Domain Prompter. Domain prompter $M_{DP}$ is designed to generate virtual domains as useful perturbations. Different from the relatively marginal improvements in 10-shot (Fig. 4(a)), our $M_{DP}$ works well for extremely low shots. The 1-shot comparison results are provided in Fig. 4(d), showing our $M_{DP}$ significantly improves the results on targets except for the UODD. Note that we don’t apply $M_{DP}$ on DeepFish since it only contains one object category. Besides quantified results, in Fig. 4(e), we provide the t-SNE features of our learned domains and the perturbed features by adding domains to the class prototypes. The 1-shot results on Clipart1K are visualized. The results reveal that: 1) the learned domains exhibit diversity; 2) adding domains to the class prototypes generates new data but keeps the semantic concepts maintained. Finally, we term these learnable perturbations as "domain" prompters rather than "style" prompters since learning meaningful styles without explicit style information is challenging. The term ’domain’ is used broadly to encompass factors, including style, that could induce changes in data distribution.
@@ 7 Implementation Details
Recall that we implement five methods for CD-FSOD as in Sec. 5.1, including “ViTDeT-FT” adapted from ViTDeT [27], Detic [53], “Detic-FT” adapted from Detic, DE-ViT [49], and our “CD-ViTO”. For all the models, as stated in Sec. 3.1, we follow the “pretrain, finetuning, and testing" pipeline, and the models pretrained on COCO are taken directly as the result of the pretrain stage. Details of how the finetuning and testing stages are implemented are as follows.
@@
Based on vanilla DE-ViT, we develop our CD-ViTO as in Sec. 4. The hyperparameter $N_{bg}$, $\alpha$ ratio of instance reweighting module, $\tau$ temperature for $\mathcal{L}_{proto}$, $\tau$ temperature for $\mathcal{L}_{domain}$ are set as 530, 0.7, 2, 0.1 for all the target datasets. While the value $N_{dom}$ means the number of virtual domains depending on the number of target classes $N$, specifically, we have $N_{dom}=2*N$. The hyperparameter Top-K ($K$) in DE-ViT is set as 5. For datasets with number of classes $N$ less than 5, we have $K=N$. We tune the trainable parameters on 1-shot around 80 epochs, and on 5/10-shot around 40 epochs. The SGD with the learning rate as 0.002 is used as the optimizer. Experiments are performed on four RTX3090 GPUs.
@@ 8.1 Results on Standard FSOD
Results are summarized in Tab. 2, as in previous works [49, 42, 49], the nAP (mAP for target data), nAP50, and nAP75 for 10/30 shots are reported. From the results, we notice that with our CD-ViTO successfully improves the performance of the base DE-ViT on all different ViT backbones and both 10 and 30 shots. This makes us build a new SOTA for FSOD. We attribute the success primarily to 1) even sets partitioned from the same dataset may have some distribution shift issues, thus our learnable instance features and domain prompter could work; 2) with the learnable instance features and instance reweighting module, our prototypes are more likely to approach the true class centers compared to averaged fixed ones.
@@ 8.2 Ablation Results on Each Module
We highlight several points. 1) overall, finetuning two heads already could achieve impressive results compared to the base DE-ViT which indicates the effectiveness of finetuning key modules in the context of CD-FSOD. 2) Equipped with our novel modules, the performance on various target datasets is further improved steadily. 3) Finetuning the $M_{RPN}$ on the base of ours may further improve results marginally, however, it would bring much more learnable parameters and thus not adopted. 4) Results also suggest that it is hard to optimize the well-trained DINOv2 backbone with few examples. Therefore, freeing the DINO backbone is not advisable. Additionally, by comparing the number of trainable parameters (No.TrainP) of ours to that "FT two heads", it is easy to tell that our novel modules are highly efficient.
@@
Recall that we set a hyperparameter $\alpha$ for fusing the deformed prototype features $fc(F_{ins}^{att})$ and the averaged prototype features $F_{ins}^{avg}$ i.e., $\alpha fc(F_{ins}^{att})+(1-\alpha)F_{ins}^{avg}$. Thus, we conduct an ablation study on the value of $\alpha$. Results are shown in Tab. 6. The finetuning, $M_{LIF}$, and $M_{IR}$ are applied. According to the results, we set $\alpha$ as 0.7.

### v3

@@ 5 Experiments
Implementation Details. The mAPs for 1/5/10 shots are reported. Besides the vanilla OVD methods that detect novel objects directly, we adopt the "pretrain, finetune, and test" pipeline for all evaluated methods as in Sec. 3.1. In the case of our CD-ViTO, the base DE-ViT model pretrained on COCO is taken, then the $M_{DET}$, $M_{CLS}$, $M_{LIF}$, $M_{IR}$, and $M_{DP}$ are tuned on novel support set $S$ using the loss as in Eq. 1. Finetuning details e.g., hyperparameters, optimizer, learning rate, and epochs will be attached in the supplementary materials.
@@ 5.1 Evaluated Methods
Typical FSOD Methods. Several FSOD methods, including Meta-RCNN [48], TFA [45], FSCE [39], and DeFRCN [33], are included to examine whether these traditional algorithms can handle the huge domain gap posed by CD-FSOD.
@@
Open-Set FSOD/OD Methods. We study DE-ViT [52], which excels in FSOD benchmarks. Considering the straightforward nature of finetuning, we also report results using the best finetuning strategy on DE-ViT, referred to as "DE-ViT-FT" (details in supplementary materials). Besides the purely vison-based DE-ViT, we evaluate another widely known vision-language matching OVD detector, namely Detic [56], and finetune it, further resulting in "Detic-FT".
@@ 5.2 Main Results on CD-FSOD
Can open-set models directly address the challenges of CD-FSOD? Unfortunately, the results of Detic and DE-ViT clearly show that the answer is no. Despite being designed to recognize any novel category, open-set detectors still face challenges posed by domain gaps. It is noteworthy that Detic performs inference using category names and the query image, essentially adopting a zero-shot approach. Consequently, the results for 1/5/10 shots remain the same. On the other hand, DE-ViT requires support instances to construct prototypes, leading to variations in results. These findings highlight the nontrivial nature of solving the CD-FSOD by directly applying existing open-set models.
@@
Do different open-set models exhibit distinct performances? Yes. We delve into the performance of DE-ViT and Detic, two models that achieve open-set learning through entirely different approaches. DE-ViT relies solely on vision features, while Detic relies on the relationship between visual features and category text. Upon analyzing the results of DE-ViT-FT and Detic-FT, a clear trend emerges. DE-ViT-FT demonstrates a decisive advantage on ArTaxOr, Clipart1k, DIOR, and DeepFish, but falls behind Detic-FT on NEU-DET and UODD. We observe that when the Instance Boundary (IB) degree of the target dataset is slight or moderate, the vision-based DE-ViT-FT performs better. Conversely, in scenarios where boundaries are significantly indefinable, indicating strong similarity between foreground and background, methods like Detic-FT, which leverages both visual and text information, prove more competent.
How do style, inter-class variance (ICV), and indefinable boundaries (IB) impact the domain gap? We meticulously analyze the influence of style, ICV, and IB on domain gaps based on experimental results and benchmark analysis in Sec. 3.3. 1) Style, often considered strongly related to the domain, has a relatively smaller impact on CD-FSOD compared to ICV and IB. For instance, DIOR shares relatively similar ICV and IB with COCO but has a different style. Despite this, the performance of typical FSOD methods does not significantly drop on DIOR. For example, DeFRCN achieves 9.3 mAP on both COCO’s testing set and DIOR under the 1-shot setting. Similarly, the cartoon-style Clipart1k, distinct from the photorealistic COCO, also receives relatively high mAPs. 2) ICV can cause domain issues but appears manageable. ArTaxOr, with the same style and similar IB as COCO, but with fine-grained categories (arthropods) compared to the coarse-level semantic concepts in COCO. This difference in ICV leads to a significant drop in performance for typical FSOD methods on ArTaxOr, such as from 9.1 to 3.6 on 1-shot using DeFRCN. However, we observe that this negative effect can be mitigated by utilizing better vision features, as demonstrated by ViTDeT-FT, DE-ViT-FT, and our CD-ViTO. 3) The change in IB poses the greatest challenge for CD-FSOD. Nearly all types of methods struggle to achieve satisfactory results on NEU-DET and UODD when significant IB exists. Blurred boundaries make it challenging to detect correct bounding boxes, resulting in poor overall performance.
@@ 5.3 Analysis on the Proposed Modules
To demonstrate the effectiveness of our proposed components, we conduct ablation studies by incrementally adding them to the base DE-ViT. The 10-shot results are presented in Fig. 4(a). Overall, we observe that: 1) By simply finetuning top heads, the performance is significantly improved. The results reinforce the observed phenomenon and analysis from Tab. 1 that finetuning is crucial in CD-FSOD. 2) The learnable instance features further boost performance clearly, indicating that we do make the features better by optimizing them. 3) Although not as pronounced as in finetuning and $M_{LIF}$, the results still demonstrate the effectiveness of our $M_{IR}$ and $M_{DP}$ modules.
@@
More Discussions on Domain Prompter. The $M_{DP}$ is designed to generate virtual domains as useful perturbations. Different from the relatively marginal improvements in 10-shot (Fig. 4(a)), the 1-shot results on Fig. 4(d) show significant gains. Note that we don’t apply $M_{DP}$ on DeepFish since it only contains one object category and the slight drop in UODD indicates that training the new parameters introduced by $M_{DP}$ in extremely data-limited cases (3 classes for UODD) may pose challenges. Besides quantified results, in Fig. 4(e), we provide the t-SNE of our learned domains and the perturbed features by adding domains to the class prototypes. The 1-shot results on Clipart1K are visualized. The results reveal that: 1) the learned domains exhibit diversity; 2) adding domains to the class prototypes generates new data but keeps the semantic concepts maintained. Finally, we term these learnable perturbations as "domain" prompters rather than "style" prompters since learning meaningful styles without explicit style information is challenging. The term "domain" is used broadly to encompass factors, including style, that could induce changes in data distribution.
@@ 7 Implementation Details
We implement six methods for CD-FSOD as in Sec. 5.1, including “ViTDeT-FT” adapted from ViTDeT [27], Detic [56], “Detic-FT” adapted from Detic, DE-ViT [52], “DE-ViT-FT” from DE-ViT, and our “CD-ViTO”. For all the models, as stated in Sec. 3.1, we follow the “pretrain, finetuning, and testing" pipeline, and the models pretrained on COCO are taken directly as the result of the pretrain stage. The finetuning and testing details are as follows.
@@
DE-ViT-FT:
We build DE-ViT-FT by finetuning the $M_{DET}$, $M_{CLS}$ of DE-ViT (as studied in Tab. 5). All the other implementation details e.g., optimizer, and learning rate are the same as our CD-ViTO.
@@
Based on DE-ViT, we develop CD-ViTO as in Sec. 4. The hyperparameter $N_{bg}$, $\alpha$ ratio of $M_{IR}$ module, $\tau$ temperature for $\mathcal{L}_{proto}$, $\tau$ temperature for $\mathcal{L}_{domain}$ are set as 530, 0.7, 2, 0.1 for all the target datasets. While the value $N_{dom}$ means the number of virtual domains depending on the number of target classes $N$, specifically, we have $N_{dom}=2*N$. The hyperparameter Top-K ($K$) in DE-ViT is set as 5. For datasets with number of classes $N$ less than 5, we have $K=N$. We tune the trainable parameters on 1-shot around 80 epochs, and on 5/10-shot around 40 epochs. The SGD with the learning rate as 0.002 is used as the optimizer. Experiments are performed on four RTX3090 GPUs.
@@ 8.1 Results on Standard FSOD
Results are summarized in Tab. 2, as in previous works [52, 45, 52], the nAP (mAP for target data), nAP50, and nAP75 for 10/30 shots are reported. From the results, we notice that with our CD-ViTO successfully improves the performance of the base DE-ViT on all different ViT backbones and both 10 and 30 shots. This makes us build a new SOTA for FSOD. We attribute the success primarily to 1) even sets partitioned from the same dataset may have some distribution shift issues; 2) with the learnable instance features our prototypes are more likely to approach the true class centers compared to fixed ones.
@@ 8.2 Ablation Results on Each Module
Ablation on Effectiveness of Modules.
Recall that in Fig. 4(a) in the main file, we demonstrate the effectiveness of modules by adding our new proposed modules one by one. To further show the effectiveness of our modules, we further conduct ablation studies that add each of the modules separately into DE-ViT-FT. The 10-shot results are shown on Tab. 4.
Results show that: 1) Basically, all the modules work well improving the DE-ViT-FT clearly. 2) Overall, we observe that $M_{LIF}$ improves the result at most, then comes the $M_{IR}$ and $M_{DP}$. 3) With all modules added, our full CD-ViTO model achieves the best results on all target datasets. The conclusions are consistent with the main file.
Table 4: The 10-shot results of ablation study on modules.
ArTaxOr Clipart1k DIOR FISH NEU-DET UODD
DE-ViT-FT 49.2 40.8 25.6 21.3 8.8 5.4
+$M_{LIF}$ 56.3 43.0 29.3 21.7 12.3 6.5
+$M_{IR}$ 56.8 42.8 28.5 21.8 12.4 5.7
+$M_{DP}$ 50.1 42.8 26.5 22.2 9.9 5.6
+ all (CD-ViTO, ours) 60.5 44.3 30.8 22.3 12.8 7.0
@@
We highlight several points. 1) overall, finetuning two heads already could achieve impressive results compared to the base DE-ViT which indicates the effectiveness of finetuning key modules in the context of CD-FSOD. 2) Equipped with our novel modules, the performance on various target datasets is further improved steadily. 3) Finetuning the $M_{RPN}$ on the base of ours may further improve results marginally, however, it would bring much more learnable parameters and thus not adopted. 4) Results also suggest that it is hard to optimize the well-trained DINOv2 backbone with few examples. Therefore, freeing the DINO backbone is not advisable. Based on the result, we set "FT two heads" as our "DE-ViT-FT". Additionally, from the number of trainable parameters (No.TrainP), it is easy to tell that our novel modules are highly efficient.
@@
Recall that we set a hyperparameter $\alpha$ for fusing the deformed prototype features $fc(F_{pro}^{att})$ and the averaged prototype features $F_{pro}^{avg}$ i.e., $\alpha fc(F_{pro}^{att})+(1-\alpha)F_{pro}^{avg}$. Thus, we conduct an ablation study on the value of $\alpha$. Results are shown in Tab. 7. The finetuning, $M_{LIF}$, and $M_{IR}$ are applied. According to the results, we set $\alpha$ as 0.7.
@@
Time Complexity and FLOPs.
To further investigate the additional cost caused by our modules, we separately add modules on DE-ViT-FT for inference and report the time complexity and FLOPs on 10-shot as in Tab. 9. From the results, we highlight that: 1) both $M_{LIF}$ and $M_{DP}$ don’t cause any cost since $M_{LIF}$ only change the feature while $M_{DP}$ is solely used during the finetuning stage. 2) The $M_{IR}$ causes minimal time and FLOPs.
Table 9: The time complexity and FLOPs of the modules on 10-shot.
+$M_{LIF}$ + $M_{IR}$ + $M_{DP}$
Time/Query +0 (only change features) +0.0006 S +0 (only for finetuning)
FLOPs +14.1 M
@@ 8.3 More Visualization Results
8.4 More Discussion with Related Works
Though discussed in related work, since AcroFOD [9], AsyFOD [8], MoFSOD [24], and Distill-cdfsod [47] all address few-shot object detection across domains. We further clarify the differences below.
Typically, 1) AcroFOD and AsyFOD assume the source and target share the same class set, differing from our task essentially, which addresses entirely different class sets. 2) MoFSOD proposes a comprehensive benchmark but lacks a specific method. Additionally, the evaluation protocols differ: MoFSOD uses natural K-shot sampling, which samples K images per class regardless of the number of object instances per image. In contrast, we use balanced sampling, ensuring exactly K object instances per class. Our manner ensures the support instances are always few-shot and align with typical FSOD practices. 3) Distill-cdfsod is the most similar work. We inherit its three target datasets and evaluation method but introduce significant novel contributions. Compared to Distill-cdfsod, we propose new metrics, explore open-set detectors, and achieve much better results. The above comparisons are summarized in Tab. 10.
Table 10: More comparison with existing works.
Task Protocol Method Metric SOTA
AcroFOD [9], AsyFOD [8] $\mathcal{C}_{S}=\mathcal{C}_{T}$
MoFSOD [24] $\mathcal{C}_{S}\cap\mathcal{C}_{T}=\emptyset$ natural K-shot $\times$ $\times$ -
Distill-cdfsod [47] $\mathcal{C}_{S}\cap\mathcal{C}_{T}=\emptyset$ balanced K-shot Distill-cdfsod $\times$ $\times$
CD-ViTO (ours) $\mathcal{C}_{S}\cap\mathcal{C}_{T}=\emptyset$ balanced K-shot CD-ViTO ICV, IB, style